OpenStudy (rayman5766):

Which could be the measures of the three angles of an acute triangle? A. 40°, 40°, 100° B. 50°, 55°, 75° C. 30°, 30°, 120° D. 30°, 60°, 90°

OpenStudy (rayman5766):

@Data_LG2 can you tell me if I am correct

OpenStudy (anonymous):

What do you think it is?

OpenStudy (anonymous):

I will

OpenStudy (anonymous):

hints: #1 angles of acute triangles are less than 90 degrees. #2 sum of angles of a triangles always equal to 180 degrees.

OpenStudy (rayman5766):

sorry I was afk for awhile and I think my answer is D

OpenStudy (rayman5766):

@Data_LG2 sorry my answer was D I was think it was d and then I had go for 15 minuets and now I am back

OpenStudy (anonymous):

\int_1^e \ln\left(\sum_{k=0}^n a_kx^k\right)\,dx

OpenStudy (welshfella):

It's not D because one angle of D = 90 degrees which is not an acute angle